# COSINE-100U: Upgrading the COSINE-100 Experiment for Enhanced Sensitivity to Low-Mass Dark Matter Detection

### v1

@@ III.3 Internal Background and Stability Measurement
After the 241Am measurements, each crystal underwent approximately two weeks of background measurements in the sea level shield setup, as shown in Fig. 6. Due to the relatively thin layers of lead and liquid scintillator, as well as the high muon flux at sea level, we could not achieve the low-background levels of the COSINE-100 experiment at Y2L. However, we were still able to study internal $\alpha$ background and assess the stability of the encapsulation.
Figure 9: Single-hit low-energy spectra of C6 at the COSINE-100 experiment (dashed line) and the sea level simple shield using COSINE-100U encapsulation (red-line)
@@
Events coincident with the liquid scintillator with energies above 80 keV and within a 200 ns coincidence window were categorized as multiple-hit events, while all other events were categorized as single-hit events. Figure 9 shows the single-hit low-energy spectra of C6 from this measurement, using the upgraded COSINE-100U encapsulation (red-solid line) compared to the same crystal in the COSINE-100 experiment. As seen in the figure, the sea level measurement had significantly higher background rates due to reduced shielding and increased muon-related backgrounds. A peak around 33 keV was observed in this sea level measurement with the COSINE-100U encapsulation. This could be due to external contributions, with the K-shell dip of non-proportional scintillation light Lee et al. (2024a) possibly contributing to the peak. Additionally, x-rays from Ba in the PMT glass, as well as Cs and In in the photocathodes, may also contribute. The removal of the 12 mm-thick quartz layer between the crystal and PMTs in the new encapsulation may have enhanced these x-ray signals.
Figure 10: (a) Charge weighted mean decay time is shown for C6. Bulk $\alpha$ (red solid rectangle) and surface $\alpha$ (blue dashed rectangle) are indicated. (b) Energy spectra of $\alpha$ candidate events in this measurement (red-solid line) is compared with COSINE-100 setup (black-dashed line)
@@
To measure internal $\alpha$ activity, we utilized charge-weighted mean decay time to distinguish between $\alpha$ and beta/gamma events, as shown in Fig. 10 (a), where $\alpha$s form a distinct cluster with shorter decay times, clearly separated from the beta/gamma events. The bulk $\alpha$ contamination from 210Po is highlighted in the red solid box, while low-energy surface $\alpha$ contamination Adhikari et al. (2024), with energy in the 1–2 MeV range, is indicated by the blue dashed box. Bulk $\alpha$ contamination originates from impurities introduced during the crystal growing process and is expected to be consistent with the COSINE-100 measurements, decreasing over time due to the decay of 210Pb ( with a half-life of 22.3 years). Surface $\alpha$ contamination, on the other hand, may occur on the crystal surface or on the PTFE reflective sheet during the encapsulation process. We used the event rate of 1–2 MeV $\alpha$-particles as an indicator of surface contamination, as shown in Fig. 10 (b). Our careful encapsulation process has minimized surface contamination. Table 2 summarizes internal $\alpha$ background measurements of the new COSINE-100U encapsulation compared to the COSINE-100 measurements near shutdown in March 2023. The bulk $\alpha$ measurements show a clear decrease in the COSINE-100U setup, consistent with the decay of internal 210Pb, while surface $\alpha$ rates are generally lower than in COSINE-100, though some crystals show slightly higher rates. We plan to systematically study surface $\alpha$ contamination by varying surface treatment methods using sample crystals to better understand the causes of contamination.
Table 2: The summary of COSINE-100U $\alpha$ measurements.
Crystal number Bulk alpha [mBq/kg] surface alpha [nBq/cm2]
@@
1 2.59 $\pm$ 0.01 2.49 $\pm$ 0.02 33.07 $\pm$ 3.70 45.22 $\pm$ 5.99
2 1.69 $\pm$ 0.01 1.63 $\pm$ 0.02 39.61 $\pm$ 3.76 30.42 $\pm$ 10.14
3 0.63 $\pm$ 0.01 0.59 $\pm$ 0.01 71.73 $\pm$ 5.06 30.89 $\pm$ 8.57
4 0.64 $\pm$ 0.01 0.60 $\pm$ 0.01 24.27 $\pm$ 2.31 51.48 $\pm$ 5.76
5 - 1.69 $\pm$ 0.02 - 99.45 $\pm$ 11.72
6 1.52 $\pm$ 0.01 1.42 $\pm$ 0.01 120.9 $\pm$ 5.90 39.11 $\pm$ 5.83
7 1.51 $\pm$ 0.01 1.43 $\pm$ 0.01 95.35 $\pm$ 5.24 61.70 $\pm$ 7.32
8 - 1.50 $\pm$ 0.02 - 24.71 $\pm$ 4.75
To verify the stability of the assembled crystals, each crystal was monitored in the sea level test measurement facility for approximately two weeks. Specifically, the 33 keV and the 46.5 keV radiation peaks were tracked to check the stability of the encapsulation. Figure 11 (a) shows the data collected over this period, plotted in 100-hour intervals, demonstrating no noticeable shifts in the peak positions. This indicates that the crystal-PMT coupling remained robust and that no infiltration of liquid scintillator or air occurred.
@@
Figure 11: Stability of the C1 encapsulation. (a) Gain stability of sea level background measurements for approximately two-week long periods was shown with stable gain condition. (b) Measurements of 241Am source to 59.54 keV at sea level measurement (black dashed line) and Yemilab measurement (red solid line) are presented. With two-months period, no significant change of gain was observed.
@@ IV.1 Decommissioning of COSINE-100
The COSINE-100 experiment, which operated at Y2L, concluded in March 2023 in preparation for the relocation of the experimental site to Yemilab Park et al. (2024); Kim and Lee (2024) and the detector upgrade for the COSINE-100U experiment. The decommissioning of the detector was completed by October 2023, as shown in Fig. 12, and all materials were delivered to Yemilab for the installation of COSINE-100U.
Figure 12: Photos taken during the decommissioning of the COSINE-100 detector at Y2L. (a) The external shield of the COSINE-100 before decommissioning. (b) Removal of outside electronics and muon detectors, with the decommissioning of lead bricks in progress. (c) All external passive shields were removed, and the PMTs for the liquid scintillator active veto detector were detached. (d) Complete removal of the COSINE-100 shield from Y2L.
@@ IV.2 Yemilab Preparation
Yemilab is a newly constructed underground laboratory in Korea, completed in September 2022, located in Jeongseon, Gangwon Province, at a depth of 1,000 m Park et al. (2024); Kim and Lee (2024). The facility offers approximately 3,000 m2 of dedicated experimental space. The underground tunnel accommodates 17 independent experimental spaces, one of which is dedicated to the COSINE-100U experiment, as shown in Fig. 13 (a). The tunnel can be accessed via a human-riding elevator through a 600 m vertical shaft and then by electric car through a 780 m horizontal access tunnel with a 12% downward slope. The surrounding rock is primarily limestone. Ongoing radioactivity measurements of rock samples using ICP-MS and HPGe detectors show that the preliminary results are generally consistent with, or slightly lower than, those from Y2L.
Figure 13: (a) Schematic view of the underground experimental area at Yemilab and (b) The COSINE-100U fridge room to operate at –30 C.
[…]
Preliminary measurements of muon flux at Yemilab indicate a flux of 1.0×$\times 10^{-7}\mu$/cm2/s, which is four times lower than the muon flux at Y2L, measured at 3.8×$\times 10^{-7}\mu$/cm2/s Prihtiadi et al. (2018). Overall, the background environments at Yemilab are significantly better than that of Y2L, leading to reduced external radioactive background contributions.
We have prepared a warehouse-type refrigerator to serve as the COSINE-100U detector room, as shown in Fig.  13 (b). The plan is to operate the COSINE-100U experiment at –30 C to enhance light yield and improve pulse shape discrimination for nuclear recoil events Lee et al. (2022). The COSINE-100U detector room measures 4 m in width, 6 m in length, and 4 m in height, and is located at the front of the COSINE tunnel. A 10 kW cryocooler will maintain the room temperature.
@@ IV.3 Shielding Installation
Figure 14: Shield installation of COSINE-100U at Yemilab. (a) The liquid scintillator active veto detector, composed of a 1 cm thick acrylic box surrounded by 3 cm think OFC, was installed with 5-inch PMTs for readout. (b) A 20 cm thick lead layer surrounds the copper container for the liquid scintillator. (c) Inside of the acrylic box for the liquid scintillator, the C5 and C8 were installed.
Inside the COSINE-100U fridge room, we installed shielding for the experiment. Most elements of the COSINE-100 experimental shield were recycled for the COSINE-100U setup to prevent external radiation from various sources and provide an active veto for internal or external contamination Adhikari et al. (2018a). This shield consists of a 4-layer nested arrangement of components, starting from the inside: 40 cm liquid scintillator, 3 cm of copper, 20 cm of lead, and 3 cm of plastic scintillator. The liquid scintillator Adhikari et al. (2021b) and plastic scintillator Prihtiadi et al. (2018) layers actively tag radioactivity from internal contamination, external radiation, and muon events.
The COSINE-100 shield utilized a steel skeleton to support heavy elements and allow access to the inner structure with a mechanical opening system Adhikari et al. (2018a). However, this design inherently included approximately 4 tons of steel inside the lead shields. In contrast, the COSINE-100U shield does not use a steel skeleton. Instead, heavy materials, such as lead bricks, are stacked on a precisely leveled steel plate, similar to the shield used in the NEON experiment Choi et al. (2023), as shown in Fig. 14. To reinforce the top structure, 5 cm×$\times$10 cm square stainless steel pipes, each 180 cm long, will support the lead bricks. Figure 15 shows the detector geometry of the experimental setup.
We produced 2,400 L Linear Alkyl-Benzene(LAB) based liquid scintillator, following a recipe similar to that used in the COSINE-100 experiment Adhikari et al. (2018a, 2021b). The old COSINE-100 liquid scintillator will be repurposed for test measurements facilities at Yemilab.
@@ IV.4 Physics Operation Plan
For the operation of the COSINE-100U, all electronics, including preamplifiers, flash analog-to-digital converters, high voltage power supplies, and the computer server for data acquisition, will be installed in a –30C environment. While the stability of individual components had already been tested at this low temperature, integrated tests at Yemilab are planned for a few weeks. During this test, only two crystals, C5 and C8, as shown in Fig. 14 (c), will be installed, and the liquid scintillator will not yet be filled. We have just started room temperature measurements to obtain reference data for this test.
If no critical issues arise from the NaI(Tl) crystal detectors, electronics, or data acquisition system, we will proceed with the installation of the remained detector components, including the other six crystals, the liquid scintillator, and the top lead bricks and outer muon plastic scintillator panels. This entire process is expected to be completed with a month. If everything proceeds smoothly, we anticipate that the COSINE-100U experiment will begin physics operations in October 2024.
@@ V Expected Background
We have gained a precise understanding of the backgrounds in the COSINE-100 detector through Geant4-based simulations Adhikari et al. (2018d, 2021c); Yu et al. (2024a). To account for COSINE-100U-specific background contributions, we constructed detector geometry for use in the Geant4-based simulation, as shown in Fig. 15. Since the COSINE-100U experiment uses the same crystals as COSINE-100, with only minor machining and surface polishing, we expect the majority of background contributions in COSINE-100U, particularly from internal contaminants, to be very similar to those observed in the COSINE-100 experiment.
Figure 15: COSINE-100U detector geometry used in the Geant4 simulation: (a) bird’s-eye view and (b) front view. Eight NaI(Tl) crystals are supported by the acrylic table inside the liquid scintillator. Nine 5-inch PMTs in each side are attached to read photons from the liquid scintillator. The outer lead shield and plastic scintillators are also installed. To support the shielding on the top, stainless steel support pipes are installed. The entire shieling setup is placed on a precisely labeled steel base.
However, a few differences are expected due to the redesigned crystal encapsulation. The encapsulation components were replaced, and an additional inner PTFE structure was incorporated. We measured the radioactivity levels of all encapsulation components, as summarized in table 3. As we carefully selected all materials, the contamination levels of the new encapsulation materials are much lower than those of the PMT and PMT base. Based on our understanding of the COSINE-100 backgrounds Yu et al. (2024a) and the measured contamination levels of the encapsulation materials, we simulated the expected background contributions. As in COSINE-100, the PMTs remain the dominant source of external background contamination.
Table 3: Measured radioactivity contaminants in detector components inside the shielding. The radioactivities were measured using high-purity germanium detectors at Y2L; upper limits are quoted with a 90% confidence level. The PMTs and PMT bases are measured in units of mBq/unit, while the other components are measured in units of mBq/kg.
@@
The polishing of all crystal surfaces and the replacement of the Teflon lapping films may result in different surface contamination levels in the COSINE-100U crystals. Generally, we observed fewer $\alpha$ particles with partial energy deposition (1–2 MeV measured energy), which may suggest lower surface contamination. However, for this study, we conservatively assume the same surface contamination background contributions as observed in the COSINE-100 experiment.
Although the same PMTs are used, the removal of the 12 mm quartz layer could potentially increase background contributions from the PMTs to the crystals. We simulated these background contributions in the COSINE-100U geometry (Fig. 15), assuming the same contamination levels as in the COSINE-100 experiment Yu et al. (2024a). The absence of the 12 mm quartz shield may enhance the x-ray contribution from the PMTs, but this effect is primarily observed at energies above 20 keV, with no significant differences in the signal region below 6 keV, as shown in Fig. 16.
Figure 16: Expected energy spectra of the C6 crystal in the COSINE-100U setup compared with the COSINE-100 background Yu et al. (2024a). (a) Low-energy spectra from the anode readout of COSINE-100 (black solid line) compared with the expected background of the COSINE-100U setup (red solid line). The main differences in PMT contributions are separated as dashed lines. (b) High-energy spectra from the dynode readout are presented in the same manner. Although an increase in background contributions from the PMTs is expected in the COSINE-100U encapsulation, its impact on the signal region below 6 keV is negligible.
Figure 16 shows the expected background levels of the COSINE-100U C6 crystal compared with those of the COSINE-100 setup, based on the assumptions discussed above and the Geant4-based simulation. The main differences are caused by PMT x-rays resulting from the absence of the 12 mm quartz windows; however, no significant difference is observed in the low-energy signal regions.

### v2

@@ III.3 Internal Background and Stability Measurement
Following the 241Am measurements, we conducted approximately two weeks of background measurements for each crystal in the sea-level shield setup as shown in Fig. 3. For this, the 241Am source was removed, as shown in Fig. 3(b). Due to the relatively thin layers of lead and liquid scintillator, as well as the high muon flux at sea level, we could not achieve the low-background levels of the COSINE-100 experiment at Y2L. However, we were still able to study internal $\alpha$ background and assess the stability of the encapsulation.
@@
Events coincident with the liquid scintillator with energies above 80 keV and within a 200 ns coincidence window were categorized as multiple-hit events, while all other events were categorized as single-hit events. Figure 6 shows the single-hit low-energy spectra of C6 from this measurement, using the upgraded COSINE-100U encapsulation (red solid line) compared to the same crystal in the COSINE-100 experiment. As seen in the figure, the sea level measurement had significantly higher background rates due to reduced shielding and increased muon-related backgrounds. A peak around 33 keV was observed in this sea level measurement with the COSINE-100U encapsulation. This could be due to cosmogenic activation of 121mTe COSINE-100:2019rvp and external contributions, with the K-shell dip of non-proportional scintillation light COSINE-100:2024log possibly contributing to the peak. Additionally, x-rays from Ba in the PMT glass, as well as Cs and In in the photocathodes, may also contribute. The removal of the 12 mm thick quartz layer between the crystal and PMTs in the new encapsulation may have enhanced these x-ray signals.
Figure 6: Background measurement at sea level test facility. The single-hit low-energy spectra of C6 from the COSINE-100 experiment (black dashed line) and the sea-level simple shield using the COSINE-100U encapsulation (red solid line) are compared. The distinct 33 keV peak arises due to cosmogenic activation of 121mTe and the removal of the quartz window in the new encapsulation, as described in Section III.3. Meanwhile, the 47 keV peak, originating from internal 210Pb, remains unchanged.
@@
To measure internal $\alpha$ activity, we utilized charge-weighted mean decay time to distinguish between $\alpha$ and beta/gamma events, as shown in Fig. 7(a), where $\alpha$s form a distinct cluster with shorter decay times, clearly separated from the beta/gamma events. The bulk $\alpha$ contamination from 210Po is highlighted in the red solid box, while low-energy surface $\alpha$ contamination COSINE-100:2023dsf , with energy in the 1–2 MeV range, is indicated by the green dashed box. Bulk $\alpha$ contamination originates from impurities introduced during the crystal growing process and is expected to be consistent with the COSINE-100 measurements, decreasing over time due to the decay of 210Pb (with a half-life of 22.3 years). Surface $\alpha$ contamination, on the other hand, may occur on the crystal surface or on the PTFE reflective sheet during the encapsulation process. We used the event rate of 1–2 MeV $\alpha$-particles as an indicator of surface contamination, as shown in Fig. 7(b). Our careful encapsulation process has minimized surface contamination. Table 3 summarizes internal $\alpha$ background measurements of the new COSINE-100U encapsulation compared to the COSINE-100 measurements near shutdown in March 2023. The bulk $\alpha$ measurements show a clear decrease in the COSINE-100U setup, consistent with the decay of internal 210Pb, while surface $\alpha$ rates are generally lower than in COSINE-100, though some crystals show slightly higher rates. We plan to systematically study surface $\alpha$ contamination by varying surface treatment methods using sample crystals to better understand the causes of contamination.
Figure 7: $\alpha$ measurements. (a) Charge weighted mean decay time is shown for C6. Bulk $\alpha$ (red solid rectangle) and surface $\alpha$ (blue dashed rectangle) are indicated. (b) Energy spectra of $\alpha$ candidate events in this measurement (red solid line) is compared with COSINE-100 setup (black dashed line). The two peaks are caused by 210Po decays, showing two different $\alpha$ quenching effects observed in the COSINE-100 experiment COSINE-100:2023dsf .
Table 3: The summary of COSINE-100U $\alpha$ measurements. Bulk $\alpha$-particles are identified within the energy range of 2.2–3.5 MeV, whicle surface $\alpha$-particles are selected in the 1–2 MeV range, as indicated in Fig. 7.
Crystal Bulk alpha [mBq/kg] surface alpha [nBq/cm2]
@@
C1 2.59 $\pm$ 0.01 2.49 $\pm$ 0.02 33.07 $\pm$ 3.70 45.22 $\pm$ 5.99
C2 1.69 $\pm$ 0.01 1.63 $\pm$ 0.02 39.61 $\pm$ 3.76 30.42 $\pm$ 10.14
C3 0.63 $\pm$ 0.01 0.59 $\pm$ 0.01 71.73 $\pm$ 5.06 30.89 $\pm$ 8.57
C4 0.64 $\pm$ 0.01 0.60 $\pm$ 0.01 24.27 $\pm$ 2.31 51.48 $\pm$ 5.76
C5 - 1.69 $\pm$ 0.02 - 99.45 $\pm$ 11.72
C6 1.52 $\pm$ 0.01 1.42 $\pm$ 0.01 120.9 $\pm$ 5.90 39.11 $\pm$ 5.83
C7 1.51 $\pm$ 0.01 1.43 $\pm$ 0.01 95.35 $\pm$ 5.24 61.70 $\pm$ 7.32
C8 - 1.50 $\pm$ 0.02 - 24.71 $\pm$ 4.75
The stability of the assembled crystals is monitored by tracking the radiation peaks: 33 keV, as explained above, and 46.5 keV from internal 210Pb. Figure 8(a) shows the data collected over this period, plotted in 100-hour intervals, demonstrating no noticeable shifts in the peak positions. This indicates that the crystal-PMT coupling remained robust and that no infiltration of liquid scintillator or air occurred.
@@
Figure 8: Stability of the COSINE-100U encapsulation. (a) Gain stability during the sea-level background measurements over a two-week period, showing stable conditions with two internal peaks for C1. (b) The 59.54 keV peak from the 241Am source measured at both sea level (black dashed line) and Yemilab (red solid line) for C1. Over a two-month period, no significant gain changes were observed in either environment.
@@ IV.1 Decommissioning of COSINE-100
The COSINE-100 experiment, which operated at Y2L, concluded in March 2023 in preparation for the relocation of the experimental site to Yemilab Park:2024sio ; Yemilab2024 and the detector upgrade for the COSINE-100U experiment. The decommissioning of the detector was completed by October 2023, as shown in Fig. 9, and all materials were delivered to Yemilab for the installation of COSINE-100U.
Figure 9: Decommissioning of the COSINE-100. Photos were taken during the decommissioning of the COSINE-100 detector at Y2L. (a) The external shield of the COSINE-100 before decommissioning. (b) Removal of outside electronics and muon detectors, with the decommissioning of lead bricks in progress. (c) All external passive shields were removed, and the PMTs for the liquid scintillator active veto detector were detached. (d) Complete removal of the COSINE-100 shield from Y2L.
@@ IV.2 Yemilab Preparation
Yemilab is a newly constructed underground laboratory in Korea, completed in September 2022, located in Jeongseon, Gangwon Province, at a depth of 1,000 m corresponding to 2,700 m water equivalent Park:2024sio ; Yemilab2024 . The facility offers approximately 3,000 m2 of dedicated experimental space. The underground tunnel accommodates 17 independent experimental spaces, one of which is dedicated to the COSINE-100U experiment, as shown in Fig. 10(a). The tunnel can be accessed via a human-riding elevator through a 600 m vertical shaft and then by electric car through a 780 m horizontal access tunnel with a 12% downward slope. The surrounding rock is primarily limestone. Ongoing radioactivity measurements of rock samples using inductively coupled plasma mass spectrometry (ICP-MS) and high purity germanium (HPGe) detectors show that the preliminary results are generally consistent with, or slightly lower than, those from Y2L.
Figure 10: Yemilab for the COSINE-100U. (a) Schematic view of the underground experimental area at Yemilab, highlighting the COSINE-100 tunnel (red circle). (b) The COSINE-100U fridge room, designed to operate at --30 C.
The ventilation system at Yemilab efficiently maintains radon levels below 50 Bq/m3 outside of the summer season. A newly installed radon-reduced air supply system keeps radon levels below 150 Bq/m3 during the summer. The post-epoxy floor coating and air filtration system have reduced PM10 dust levels to below 10 $\mu$g/m3, well within typical office environmental standards. Stricter controls aim to further reduce dust level to below 5 $\mu$g/m3. Additionally, Yemilab features a Radon Reduction System (RRS) supplying 50 m3/h of air with radon levels below 100 mBq/m3, which will be used in the COSINE-100U detector room.
Preliminary measurements of the muon flux in Yemilab indicate a flux of 1.0×$\times 10^{-7}\mu$/cm2/s, which is four times lower than the muon flux in Y2L, measured at 3.8×$\times 10^{-7}\mu$/cm2/s Prihtiadi:2017inr . Overall, the background environments at Yemilab are significantly better than that of Y2L, leading to reduced external radioactive background contributions.
We have prepared a warehouse-type refrigerator with a 10 kW cryocooler to serve as the COSINE-100U detector room, as shown in Fig.  10(b). The plan is to operate the COSINE-100U experiment at --30 C to enhance light yield and improve pulse shape discrimination for nuclear recoil events Lee:2021aoi . Based on previous measurements at --35 C, we expect an increase in light yield of approximately 5% compared to operation at room temperature for electron recoil events. Additionally, an increase in the $\alpha$ quenching factor of approximately 9% was observed, suggesting a potential further improvement in light yield for nuclear recoil events. While cooling to --35 C is technically feasible, operating at --30 C was choosen to avoid excessive load on the cryocooler. The COSINE-100U detector room measures 4 m in width, 6 m in length, and 4 m in height, and is located at the front of the COSINE tunnel.
@@ IV.3 Shielding Installation
Inside the COSINE-100U fridge room, shielding was installed to protect the experiment from external radiation sources and to provide an active veto for internal or external contamination Adhikari:2017esn . Most of the shielding components from the COSINE-100 experiment were recycled for use in the COSINE-100U setup. This shielding consists of a four-layer nested arrangement, starting from the inside: 40 cm of liquid scintillator, 3 cm of copper, 20 cm of lead, and 3 cm of plastic scintillator. The liquid scintillator Adhikari:2020asl and plastic scintillator Prihtiadi:2017inr layers actively tag radioactivity from internal contamination, external radiation, and muon events.
The COSINE-100 shield utilized a steel skeleton to support heavy elements and allow access to the inner structure with a mechanical opening system Adhikari:2017esn . However, this design inherently included approximately 4 tons of steel inside the lead shields. In contrast, the COSINE-100U shield does not use a steel skeleton. Instead, heavy materials, such as lead bricks, are stacked on a precisely leveled steel plate, similar to the shield used in the NEON experiment NEON:2022hbk , as shown in Figs. 11 and 12. To reinforce the top structure, 5 cm×$\times$10 cm square stainless steel pipes, each 180 cm in length, will support the lead bricks. Figure 11 illustrates the overall detector geometry of the COSINE-100U setup, while Fig. 12 shows a photograph of the COSINE-100U shield during installation at Yemilab.
Figure 11: Schematic of the COSINE-100U detector. The setup includes eight NaI(Tl) detectors with a total mass of 99.1 kg, immersed in 2,200 L of liquid scintillator. The detectors are surrounded by multiple layers of shielding: 3 cm of copper, 20 cm of lead, and 3 cm of plastic scintillator, effectively minimizing background contributions.
We produced 2,400 L Linear Alkyl-Benzene(LAB) based liquid scintillator Kim:2024spf , following a recipe similar to that used in the COSINE-100 experiment Adhikari:2017esn ; Adhikari:2020asl . The old COSINE-100 liquid scintillator will be repurposed for test measurements facilities at Yemilab.
Figure 12: Shield installation of COSINE-100U at Yemilab. (a) The liquid scintillator active veto detector, composed of a 1 cm thick acrylic box surrounded by 3 cm think OFC, was installed with 5-inch PMTs for readout. (b) A 20 cm thick lead layer surrounds the copper container for the liquid scintillator. (c) Inside of the acrylic box for the liquid scintillator, the C5 and C8 crystals were installed.
@@ IV.4 Physics Operation Plan
For the operation of COSINE-100U, all electronics–including preamplifiers, FADCs, high-voltage power supplies, and the computer server for data acquisition–will be installed in a --30C environment. We have tested all these electronics in a --30C environment for three months, and no issues were observed.
Increased light yield and improved pulse shape discrimination of nuclear recoil events for NaI(Tl) crystals at --30C were observed, as reported in Ref. Lee:2021aoi . Preliminary tests with a small liquid scintillator cell also showed an increased light yield, consistent with results from the literature Xia:2014cca .
However, initial measurements of two crystals (C5 and C8) installed at Yemilab under --30C operation, as shown in Fig. 12(c), revealed weakening of the sealing through the PTFE gasket. This issue may be due to differential thermal contraction between the PTFE gasket and the OFC lid. To address this, the PTFE gasket is being replaced with a Viton O-ring to ensure reliable sealing at low temperatures.
We assembled a test crystal from the same manufacturer as the COSINE-100 crystals using the Viton O-ring to assess stability at room temperature. During two months of measurements, no performance issues were observed. The modification of the gaskets was applied to all COSINE-100U crystals.
Once all detectors are assembled, we will proceed with the installation of all COSINE-100U components, including the eight crystals, the liquid scintillator, and the top lead bricks and outer muon plastic scintillator panels. If the schedule proceeds as planned, the COSINE-100U experiment will begin physics operations at room temperature in 2025.
Simultaneously, the test crystal will undergo stability checks at low temperature using a refrigerator installed at IBS. A test period of at least six months is planned to confirm its long-term stability. Once sufficient confidence in low-temperature operation is achieved, the experiment will transit to low-temperature operation.
@@ V Expected Background
We have gained a precise understanding of the backgrounds in the COSINE-100 detector through Geant4-based simulations cosinebg ; cosinebg2 ; COSINE-100:2024ola . To account for COSINE-100U-specific background contributions, we constructed detector geometry for use in the Geant4-based simulation, as shown in Fig. 13. Since the COSINE-100U experiment uses the same crystals as COSINE-100, with only minor machining and surface polishing, we expect the majority of background contributions in COSINE-100U, particularly from internal contaminants, to be very similar to those observed in the COSINE-100 experiment.
Figure 13: COSINE-100U detector geometry used in the Geant4 simulation. (a) Bird’s-eye view and (b) front view. Eight NaI(Tl) crystals are supported by the acrylic table inside the liquid scintillator. Nine 5-inch PMTs in each side are attached to read photons from the liquid scintillator. The outer lead shield and plastic scintillators are also installed. To support the shielding on the top, stainless steel support pipes are installed. The entire shielding setup is placed on a precisely labeled steel base.
However, a few differences are expected due to the redesigned crystal encapsulation. The encapsulation components were replaced, and an additional inner PTFE structure was incorporated. We measured the radioactivity levels of all encapsulation components, as summarized in Table 4. As we carefully selected all materials, the contamination levels of the new encapsulation materials are much lower than those of the PMT and PMT base. Based on our understanding of the COSINE-100 backgrounds COSINE-100:2024ola and the measured contamination levels of the encapsulation materials, we simulated the expected background contributions of radiation components. As in COSINE-100, the PMTs remain the dominant source of external background contamination.
Table 4: Measured radioactivity contaminants in detector components inside the shielding. The radioactivities were measured using high-purity germanium detectors at Y2L assuming secular equilibrium; upper limits are quoted with a 90% confidence level. The PMTs, PMT bases, and Viton O-ring are measured in units of mBq/unit, while the other components are measured in units of mBq/kg.
@@
Viton O-ring 3.5 ± 0.4 0.8 ± 0.1 4.8 ± 1.2
The polishing of all crystal surfaces and the replacement of the Teflon lapping films may result in different surface contamination levels in the COSINE-100U crystals. Generally, we observed fewer $\alpha$ particles with partial energy deposition (1–2 MeV measured energy), which may suggest lower surface contamination (see Section III.3). However, for this study, we conservatively assume the same surface contamination background contributions as observed in the COSINE-100 experiment.
Although the same PMTs are used, the removal of the 12 mm quartz layer could potentially increase background contributions from the PMTs to the crystals. We simulated these background contributions in the COSINE-100U geometry (Fig. 13), assuming the same contamination levels as in the COSINE-100 experiment COSINE-100:2024ola . The absence of the 12 mm quartz shield may enhance the x-ray contribution from the PMTs, but this effect is primarily observed at energies above 20 keV, with no significant differences in the signal region below 6 keV, as shown in Fig. 14.
Figure 14: Expected energy spectra of the C6 crystal. The expected background level of the C6 in the COSINE-100U setup compared with the COSINE-100 background COSINE-100:2024ola . (a) Low-energy spectra from the anode readout of COSINE-100 (black solid line) compared with the expected background of the COSINE-100U setup (red solid line). The main differences in PMT contributions are separated as dashed lines. (b) High-energy spectra from the dynode readout are presented in the same manner. Although an increase in background contributions from the PMTs is expected in the COSINE-100U encapsulation, its impact on the signal region below 6 keV is negligible.